OpenStudy (mvpriest6):

Match each set of ordered pairs with their rise/run terms: (2,-3) and (-1,4), (-2,3) and (-1,4), (2,-3) and (1,4), (-2,3) and (1,4) definitions: a) 1 b) -7/3 c)-7 d)1/3

OpenStudy (mvpriest6):

@iGreen can you plz help me

OpenStudy (mvpriest6):

 isnt it that type of thing

OpenStudy (igreen):

Ah, it wants you to match the points with their slopes.

OpenStudy (mvpriest6):

so i was right

OpenStudy (igreen):

We have 4 sets of ordered pairs: (2,-3) and (-1,4) (-2,3) and (-1,4) (2,-3) and (1,4) (-2,3) and (1,4) We plug these into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\). \((2, -3), (-1, 4)\) \(m = \dfrac{4+3}{-1-2}\) \((-2, 3), (-1, 4)\) \(m = \dfrac{4 - 3}{-1+2}\) \((2, -3), (1, 4)\) \(m = \dfrac{4 + 3}{1-2}\) \((-2, 3), (1, 4)\) \(m = \dfrac{4 - 3}{1+2}\) Can you simplify and solve all four of those? @Mvpriest6

OpenStudy (mvpriest6):

i can do it thx

OpenStudy (igreen):

No problem, tell me what you get.

OpenStudy (mvpriest6):

ok

OpenStudy (mvpriest6):

first one is 1/3

OpenStudy (igreen):

Um, how did you get that?

OpenStudy (mvpriest6):